Please help me solve this word problem as I don't understand how to solve mixture problems.
What quantity of 70 per cent acid solution must be mixed with a 25 solution to produce 972 mL of a 50 per cent solution?

Found 2 solutions by lwsshak3, rothauserc:
Answer by lwsshak3(11628)

Please help me solve this word problem as I don't understand how to solve mixture problems.
What quantity of 70 per cent acid solution must be mixed with a 25 solution to produce 972 mL of a 50 per cent solution?
***
let x=quantity (mL) of 70% solution to be mixed
972-x=quantity (mL) of 25% solution to be mixed
quantity of acid in the two solutions=quantity of acid in the mixture
70%x+25%(972-x)=50%*972
.70x+243-.25x=486
.45x=243
x=540
972-x=432
quantity (mL) of 70% solution to be mixed=540 mL
quantity (mL) of 25% solution to be mixed=432 mL

Answer by rothauserc(4718)

First multiply 972 * .50 = 486 ml of acid because it is a 50% solution. Now,
.70x + .25x = 486
.95x = 486
x = 486 / .95 = 511.58 and
.70 * 511.58 = 358.11 mL of the 70% solution